# Styx: Transactional Stateful Functions on Streaming Dataflows

### v2

@@ 8.3. Scalability
In this experiment, we tested the scalability of Styx by increasing the number of Styx workers (1 CPU each). We measure the maximum throughput on YCSB-T. The goal is to calculate the speedup of operations as the input throughput and number of workers scale together. In addition, we set a parameter to control the percentage of multi-partition transactions in the workload. In Figure 12, we observe that in all settings, Styx retains near-linear scalability. The percentage of multi-partition transactions in the workload does not affect performance significantly, albeit YCSB-T consists of two key accesses per transaction.
8.4. Snapshotting Microbenchmarks
@@
Effect of Snapshots. In Figure 13 we depict the impact of the asynchronous incremental snapshots to Styx’s performance. In both figures, we display when a snapshot starts and ends. The state includes 1 million keys with a 1-second snapshot interval. Styx is deployed with four 1-CPU workers, and the input transaction arrival rate is fixed to 3K YCSB-T TPS. In Figure 12(a), we observe that during a snapshot operation, Styx shows no performance degradation in the output throughput as it closely follows the input. In Figure 12(b), we observe a minor increase in the end-to-end latency in some snapshots. This is normal, as the concurrent snapshotting thread competes with the transaction execution thread during snapshotting, and it also has to block the transaction execution thread momentarily in order to copy the corresponding operator’s delta state.
@@
Recovery Time. In Figure 14, we evaluate the recovery process of Styx with the same parameters as in Figure 13. We reboot a Styx worker at ~13.5 seconds. It takes Styx’s coordinator roughly a second to detect the failure. Then, after the reboot, the coordinator re-registers the worker and notifies all workers to load the last complete snapshot, merge any uncompacted deltas, and use the message broker offsets of that snapshot. The recovery time is also observed in the latency (Figure 13(b)) that is ~2.5 seconds (time to detect the failure + time to complete recovery). In terms of throughput (Figure 13(a)), we observe Styx working on its maximum throughput after recovery completes to keep up with the backlog and the input throughput.
@@
Effect of Large State Snapshots. In Figure 15, we test the snapshotting mechanism against a larger state of 20 GB from TPC-C using a bigger Styx deployment of 100 1-CPU workers at 10-second checkpoint intervals. From 0 to the 750-second mark, Styx is importing the dataset. Since there are no small deltas (importing is an append-only operation), snapshotting is more expensive than the normal workload execution, where only the deltas are stored in the snapshots. The rise in latency at ~550 seconds corresponds to the loading of the largest tables (Stock and Order-Line) in the system. After loading the data and starting the transactional workload at 1000 TPS, we observe a drop in latency due to fewer state changes within the delta maps.
@@ 9. Related Work
Transactional SFaaS. SFaaS has received considerable research attention and open-source work. Transactional support alongside fault tolerance guarantees (that popularized DBMS systems) is necessary to widen the adoption of SFaaS. Existing systems fall into two categories: i) those that focus on transactional serializability and ii) those that provide eventual consistency. The first category includes Beldi (Zhang et al., 2020), Boki (Jia and Witchel, 2021a), and T-Statefun (de Heus et al., 2021a). Beldi implements linked distributed atomic affinity logging on DynamoDB to guarantee serializable transactions among AWS Lambda functions with a variant of the two-phase commit protocol. Boki extends Beldi by adding transaction pipeline improvements regarding the locking mechanism and workflow re-execution. T-Statefun (de Heus et al., 2021a) also uses two-phase commit with coordinator functions to support serializability on top of Apache Flink’s Statefun. For eventually consistent transactions, T-Statefun provides an alternative option to two-phase commit using the Sagas pattern. Cloudburst (Sreekanti et al., 2020) also provides causal consistency guarantees within a DAG workflow. Proposed more recently, Netherite (Burckhardt et al., 2022) offers exactly-once guarantees and a high-level programming model for Microsoft’s Durable Functions (Burckhardt et al., 2021b), but it does not guarantee transactional serializability across functions.
Serverless Runtimes. Serverless runtimes are an orthogonal line of research to transactional SFaaS. They support execution in a sandboxed runtime environment to make (stateless) serverless execution efficient, performant, secure, and developer-friendly. Faasm (Shillaker and Pietzuch, 2020), and Shredder (Zhang et al., 2019) provide low-latency implementations of SFaaS with WebAssembly support. Nightcore (Jia and Witchel, 2021c) is another runtime that utilizes threads in long-running containers, used by Boki (Jia and Witchel, 2021a).
@@
Dataflow Systems. Support for fault-tolerant execution in the cloud with exactly-once guarantees (Fernandez et al., 2014; Carbone et al., 2017) is one of the main drivers behind the wide adoption of modern dataflow systems. However, they lack a generic and developer-friendly programming model with support for transactions and a natural way to program function-to-function calls. Closer to the spirit of Styx are Ciel (Murray et al., 2011) and Noria (Gjengset et al., 2018). Ciel proposes a language and an engine for distributed fault-tolerant computations based on functions. At the same time, Noria solves the view maintenance problem via a dataflow architecture that can propagate updates to clients quickly, targeting web-based, read-heavy computations. Yet, none of the two provide a serializable transactional model for workflows of functions like Styx.
Elasticity in Dataflow Systems. A lot of work has been carried out in dynamic reconfiguration (Gjengset et al., 2018; Kalavri et al., 2018; Floratou et al., 2017) and state migration (Hoffmann et al., 2019; Gu et al., 2022; Del Monte et al., 2020) of streaming dataflow systems over the last few years. These advancements are necessary for providing serverless elasticity in the case of state and compute collocation and enable the leveraging of dataflow graphs as an execution model for serverless stateful cloud applications, which is a future goal of Styx.
@@
Transactional Protocols. Besides Aria (Lu et al., 2020) that inspired the protocol we created for Styx Section 4, two other protocols fit the requirement of no a-priori read/write set knowledge: Starry (Zhang et al., 2022) and Lotus (Zhou et al., 2022). Starry targets replicated databases with a semi-leader protocol for multi-master transaction processing. At the same time, Lotus (Zhou et al., 2022) focuses on improving the performance of multi-partition workloads using a new methodology called run-to-completion-single-thread (RCST). Styx makes orthogonal contributions to these works and could adopt multiple ideas from them in the future.
Benchmarks. Due to the absence of benchmarks for (transactional) SFaaS, we used an extended variant of YCSB (Cooper et al., 2010) inspired by T-Statefun’s (de Heus et al., 2021a) evaluation. Another benchmark we employ is a variant of Deathstar (Gan et al., 2019), adapted to SFaaS workloads by (Zhang et al., 2020). However, it contains a small percentage of transactions (0.5%). Finally, a transactional benchmark that targets OLTP systems is TPC-C (Leutenegger and Dias, 1993), and to use it in our evaluation, we had to rewrite it for the SFaaS paradigm.
Summary. To the best of our knowledge, Styx is the first SFaaS system that performs deterministic transactions, leverages a dataflow engine that encapsulates both state and function execution, and provides a high-level SFaaS programming model that hides low-level transaction processing primitives from programmers. Finally, Styx outperforms the SotA by more than an order of magnitude in terms of throughput.
@@ Acknowledgements.
This publication is part of project number 19708 of the Vidi research program, which is partly financed by the Dutch Research Council (NWO).References
@@
- Cooper et al. (2010) Brian F Cooper, Adam Silberstein, Erwin Tam, Raghu Ramakrishnan, and Russell Sears. 2010. Benchmarking cloud serving systems with YCSB. In Proceedings of the 1st ACM symposium on Cloud computing. 143–154.
- de Heus et al. (2021a) Martijn de Heus, Kyriakos Psarakis, Marios Fragkoulis, and Asterios Katsifodimos. 2021a. Distributed transactions on serverless stateful functions. In Proceedings of the 15th ACM International Conference on Distributed and Event-based Systems. 31–42.
- de Heus et al. (2021b) Martijn de Heus, Kyriakos Psarakis, Marios Fragkoulis, and Asterios Katsifodimos. 2021b. Distributed transactions on serverless stateful functions. In DEBS.
@@
- Jia and Witchel (2021c) Zhipeng Jia and Emmett Witchel. 2021c. Nightcore: efficient and scalable serverless computing for latency-sensitive, interactive microservices. In Proceedings of the 26th ACM International Conference on Architectural Support for Programming Languages and Operating Systems. 152–166.
@@
- Lamport (2001) Leslie Lamport. 2001. Paxos made simple. ACM SIGACT News (Distributed Computing Column) 32, 4 (Whole Number 121, December 2001) (2001), 51–58.
@@
- Ongaro and Ousterhout (2014) Diego Ongaro and John Ousterhout. 2014. In search of an understandable consensus algorithm. In ATC.
- Shillaker and Pietzuch (2020) Simon Shillaker and Peter Pietzuch. 2020. Faasm: Lightweight isolation for efficient stateful serverless computing. In 2020 USENIX Annual Technical Conference (USENIX ATC 20). 419–433.
@@
- Zhang et al. (2019) Tian Zhang, Dong Xie, Feifei Li, and Ryan Stutsman. 2019. Narrowing the gap between serverless and its state with storage functions. In Proceedings of the ACM Symposium on Cloud Computing. 1–12.

### v4

@@ 8.3. Scalability
In this experiment, we test the scalability of Styx by increasing the number of Styx workers. Each worker is assigned 1 CPU each and a state of 1 million keys. We measure the maximum throughput on YCSB-T. The goal is to calculate the speedup of operations as the input throughput and number of workers scale together. In addition, we control the percentage of multi-partition transactions in the workload, i.e., transactions that span across workers. In Figure 12, we observe that in all settings, Styx retains near-linear scalability. Finally, Styx displays the expected behavior as the number of multi-partition transactions increases.
8.4. Fault-Tolerance Evaluation
@@
Effect of Snapshots. In Figure 13, we depict the impact of the asynchronous incremental snapshots to Styx’s performance. In both figures, we mark when a snapshot starts and ends. The state includes 1 million keys, and we use a 1-second snapshot interval. Styx is deployed with four 1-CPU workers, and the input transaction arrival rate is fixed to 3K YCSB-T TPS. In Figure 13(a), we observe that during a snapshot operation, Styx shows virtually no performance degradation in throughput. In Figure 13(b), we observe a minor increase in the end-to-end latency in some snapshots. The reason for that is the concurrent snapshotting thread, which competes with the transaction execution thread during snapshotting. At the same time, it also has to block the transaction execution thread momentarily to copy the corresponding operator’s state delta.
@@
Recovery Time. In Figure 14, we evaluate the recovery process of Styx with the same parameters as in Figure 13. We reboot a Styx worker at ~13.5 seconds. It takes Styx’s coordinator roughly a second to detect the failure. Then, after the reboot, the coordinator re-registers the worker and notifies all workers to load the last complete snapshot, merge any uncompacted deltas, and use the message broker offsets of that snapshot. The recovery time is also observed in the latency (Figure 14(b)) that is ~2.5 seconds (time to detect the failure in addition to the time to complete recovery). In terms of throughput (Figure 14(a)), we observe Styx working on its maximum throughput after recovery completes to keep up with the backlog and the input throughput.
@@
Effect of Large State Snapshots. In Figure 15, we test the incremental snapshotting mechanism against a larger state of 20 GB from TPC-C using a bigger Styx deployment of 100 1-CPU workers at 10-second checkpoint intervals. From 0 to the 750-second mark, Styx is importing the dataset. Since there are no small deltas (importing is an append-only operation), snapshotting is more expensive than the normal workload execution, where only the deltas are stored in the snapshots. The increase in latency at ~550 seconds corresponds to the loading of the largest tables (Stock and Order-Line) in the system. After loading the data and starting the transactional workload at 1000 TPS, we observe a drop in latency due to fewer state changes within the delta maps.
@@ 9. Related Work
Transactional SFaaS. SFaaS has received considerable research attention and open-source work. Transactional support with fault tolerance guarantees (that popularized DBMS systems) is necessary to widen the adoption of SFaaS. Existing systems fall into two categories: i) those that focus on transactional serializability and ii) those that provide eventual consistency. The first category includes Beldi (Zhang et al., 2020), Boki (Jia and Witchel, 2021a), and T-Statefun (de Heus et al., 2021). Beldi implements linked distributed atomic affinity logging on DynamoDB to guarantee serializable transactions among AWS Lambda functions with a variant of the two-phase commit protocol. Boki extends Beldi by adding transaction pipeline improvements regarding the locking mechanism and workflow re-execution. In turn, Halfmoon (Qi et al., 2023) extends Boki with an optimal logging implementation. T-Statefun (de Heus et al., 2021) also uses two-phase commit with coordinator functions to support serializability on top of Apache Flink’s Statefun. For eventually consistent transactions, T-Statefun implements the Sagas pattern. Cloudburst (Sreekanti et al., 2020) also provides causal consistency guarantees within a DAG workflow. Proposed more recently, Netherite (Burckhardt et al., 2022) offers exactly-once guarantees and a high-level programming model for Microsoft’s Durable Functions (Burckhardt et al., 2021b), but it does not guarantee transactional serializability across functions. Unum (Liu et al., 2023) needs to be paired with Beldi or Boki to ensure end-to-end exactly-once and transactional guarantees.
@@
Dataflow Systems. Support for fault-tolerant execution in the cloud with exactly-once guarantees (Fernandez et al., 2014; Carbone et al., 2017) is one of the main drivers behind the wide adoption of modern dataflow systems. However, they lack a general and developer-friendly programming model with support for transactions and a natural way to program function-to-function calls. Closer to the spirit of Styx are Ciel (Murray et al., 2011) and Noria (Gjengset et al., 2018). Ciel proposes a language and runtime for distributed fault-tolerant computations that can execute control flow. Noria solves the view maintenance problem via a dataflow architecture that can propagate updates to clients quickly, targeting web-based, read-heavy computations. However, none of the two provide a transactional model for workflows of functions like Styx.
@@
Transactional Protocols. Besides Aria (Lu et al., 2020) that inspired the protocol we created for Styx Section 4, two other protocols fit the requirement of no a priori read/write set knowledge: Starry (Zhang et al., 2022) and Lotus (Zhou et al., 2022). Starry targets replicated databases with a semi-leader protocol for multi-master transaction processing. At the same time, Lotus (Zhou et al., 2022) focuses on improving the performance of multi-partition workloads using a new methodology called run-to-completion-single-thread (RCST). Styx makes orthogonal contributions to these works and could adopt multiple ideas from them in the future.
10. Future Work
Elasticity in Dataflow Systems. Extensive work has been carried out in dynamic reconfiguration (Gjengset et al., 2018; Kalavri et al., 2018; Floratou et al., 2017) and state migration (Hoffmann et al., 2019; Gu et al., 2022; Del Monte et al., 2020) of streaming dataflow systems over the last few years. These advancements are necessary for providing serverless elasticity in the case of state and compute collocation to leverage dataflows as an execution model for serverless stateful cloud applications, which is a future goal of Styx.
Replication for High Availability. In the Styx architecture, replication is only applied in the snapshot store and the Input/Output queues to ensure fault tolerance. For high-availability, Styx could adopt replication mechanisms from deterministic databases. Specifically, the design of deterministic transaction protocols, such as Calvin (Thomson et al., 2012), feature state replicas that require no explicit synchronization. First, the sequencer replicas need to agree on the order of execution. After that, the deterministic sequencing algorithm guarantees that the resulting state will be the same across partition/worker replicas by all replicas executing state updates in the same order.
Non-Deterministic Functions on Streaming Dataflows. In its current version, Styx requires application logic to be deterministic, similar to OLTP (Stonebraker and Weisberg, 2013; Kallman et al., 2008), where stored procedures are required to be deterministic since they run independently on different replicas. The same determinism requirement applies to SFaaS (de Heus et al., 2021; Jia and Witchel, 2021a) systems. However, real-world applications may encapsulate logic that makes the outcome of their execution non-deterministic. Examples of non-deterministic operations are calls to external systems and using random number generators or time-related activities. That said, we have a plan for supporting non-deterministic functions in Styx, as discussed in Section 7.5.
@@ Acknowledgements.
We want to thank Paris Carbone for his advice throughout the process of developing Styx and the anonymous reviewers for their constructive feedback. This publication is part of project number 19708 of the Vidi research program, partly financed by the Dutch Research Council (NWO).References
@@
- Beillahi et al. (2022) Sidi Mohamed Beillahi, Ahmed Bouajjani, Constantin Enea, and Shuvendu Lahiri. 2022. Automated Synthesis of Asynchronizations. In International Static Analysis Symposium. Springer, 135–159.
@@
- Cheung et al. (2021) Alvin Cheung, Natacha Crooks, Joseph M. Hellerstein, and Matthew Milano. 2021. New Directions in Cloud Programming. In CIDR.
- Corporation (2025) Microsoft Corporation. 2025. Azure Logic Apps. https://azure.microsoft.com/en-us/products/logic-apps/. Accessed: 14-01-2025.
- de Heus et al. (2021) Martijn de Heus, Kyriakos Psarakis, Marios Fragkoulis, and Asterios Katsifodimos. 2021. Distributed transactions on serverless stateful functions. In Proceedings of the 15th ACM International Conference on Distributed and Event-based Systems. 31–42.
@@
- Fragkoulis et al. (2024) Marios Fragkoulis, Paris Carbone, Vasiliki Kalavri, and Asterios Katsifodimos. 2024. A survey on the evolution of stream processing systems. The VLDB Journal 33, 2 (2024), 507–541.
@@
- IETF (2025) IETF. 2025. The Idempotency-Key HTTP Header Field. https://www.ietf.org/archive/id/draft-ietf-httpapi-idempotency-key-header-05.txt. Accessed: 14-01-2025.
@@
- Kallman et al. (2008) Robert Kallman, Hideaki Kimura, Jonathan Natkins, Andrew Pavlo, Alexander Rasin, Stanley Zdonik, Evan P. C. Jones, Samuel Madden, Michael Stonebraker, Yang Zhang, John Hugg, and Daniel J. Abadi. 2008. H-Store: A High-Performance, Distributed Main Memory Transaction Processing System. In VLDB.
@@
- Kreps et al. (2011) Jay Kreps, Neha Narkhede, Jun Rao, et al. 2011. Kafka: A distributed messaging system for log processing. In Proceedings of the NetDB, Vol. 11. 1–7.
@@
- Li et al. (2023) Tianyu Li, Badrish Chandramouli, Sebastian Burckhardt, and Samuel Madden. 2023. DARQ Matter Binds Everything: Performant and Composable Cloud Programming via Resilient Steps. Proc. ACM Manag. Data 1, 2, Article 117 (June 2023), 27 pages. https://doi.org/10.1145/3589262
- Li et al. (2024) Tianyu Li, Badrish Chandramouli, Sebastian Burckhardt, and Samuel Madden. 2024. Serverless State Management Systems.. In CIDR.
- Liu et al. (2023) David H. Liu, Amit Levy, Shadi A. Noghabi, and Sebastian Burckhardt. 2023. Doing More with Less: Orchestrating Serverless Applications without an Orchestrator. In 20th USENIX Symposium on Networked Systems Design and Implementation (NSDI 23). 1505–1519.
@@
- Psarakis et al. (2025) Kyriakos Psarakis, George Christodoulou, Marios Fragkoulis, and Asterios Katsifodimos. 2025. Transactional Cloud Applications Go with the (Data)Flow. In 15th Annual Conference on Innovative Data Systems Research (CIDR’25). January 19-22, 2025, Amsterdam, The Netherlands.
- Psarakis et al. (2023) Kyriakos Psarakis, Wouter Zorgdrager, Marios Fragkoulis, Guido Salvaneschi, and Asterios Katsifodimos. 2023. Stateful Entities: Object-oriented Cloud Applications as Distributed Dataflows. In 13th Conference on Innovative Data Systems Research, CIDR 2023, Amsterdam, The Netherlands, January 8-11, 2023. www.cidrdb.org. https://www.cidrdb.org/cidr2023/papers/p70-katsifodimos.pdf
- Qi et al. (2023) Sheng Qi, Xuanzhe Liu, and Xin Jin. 2023. Halfmoon: Log-optimal fault-tolerant stateful serverless computing. In Proceedings of the 29th Symposium on Operating Systems Principles. 314–330.
- Siachamis et al. (2024) George Siachamis, Kyriakos Psarakis, Marios Fragkoulis, Arie van Deursen, Paris Carbone, and Asterios Katsifodimos. 2024. CheckMate: Evaluating Checkpointing Protocols for Streaming Dataflows. In 2024 IEEE 40th International Conference on Data Engineering (ICDE). 4030–4043. https://doi.org/10.1109/ICDE60146.2024.00309
@@
- Skiadopoulos et al. (2021) Athinagoras Skiadopoulos, Qian Li, Peter Kraft, Kostis Kaffes, Daniel Hong, Shana Mathew, David Bestor, Michael Cafarella, Vijay Gadepally, Goetz Graefe, et al. 2021. DBOS: a DBMS-oriented Operating System. (2021).
- Spenger et al. (2022) Jonas Spenger, Paris Carbone, and Philipp Haller. 2022. Portals: an extension of dataflow streaming for stateful serverless. In Proceedings of the 2022 ACM SIGPLAN International Symposium on New Ideas, New Paradigms, and Reflections on Programming and Software. 153–171.
@@
- Stonebraker et al. (2005) Michael Stonebraker, Uǧur Çetintemel, and Stan Zdonik. 2005. The 8 requirements of real-time stream processing. ACM Sigmod Record 34, 4 (2005), 42–47.
- Stonebraker and Weisberg (2013) Michael Stonebraker and Ariel Weisberg. 2013. The VoltDB Main Memory DBMS. IEEE Data Eng. Bull. 36, 2 (2013), 21–27.
- Temporal (2025) Temporal. 2025. Introducing Temporal .NET – Deterministic Workflow Authoring in .NET. https://temporal.io/blog/introducing-temporal-dotnet. Accessed: 14-01-2025.
@@
- Zhang et al. (2024) Shuhao Zhang, Juan Soto, and Volker Markl. 2024. A survey on transactional stream processing. The VLDB Journal 33, 2 (2024), 451–479.